OpenStudy (anonymous):

48 is 35% of what number?

OpenStudy (anonymous):

@kaylalynn and @shelbygt520

OpenStudy (anonymous):

same thing we did earlier

OpenStudy (anonymous):

is it 14.4?

OpenStudy (anonymous):

Say?

OpenStudy (anonymous):

\[\frac{ 48 }{ x } = \frac{ 35 }{ 100 }\] \[48*100=4800/35x\] \[=137.14\]

OpenStudy (anonymous):

because this time we know the number in which = 35% of another number, we do not know the number of which 48 is 35%